OpenStudy (anonymous):

Find the slope of the line that passes through the points (2,-3) and (-5,1)

jimthompson5910 (jim_thompson5910):

Remember that the two generic points are (x1, y1) and (x2, y2) (x1, y1) = (2, -3), which means x1=2 and y1=-3 and (x2, y2) = (-5, 1), which means x2=-5 and y2=1 So x1=2, y1=-3, x2=-5, and y2=1. ------------------- Now use the slope formula m = (y2-y1)/(x2-x1) m = (1-(-3))/(-5-2) m = (1+3)/(-5-2) m = 4/(-7) m = -4/7 So the slope of the line that passes through the two points is -4/7

OpenStudy (amistre64):

(2,-3) -(-5,1) ------- <7,-4>; slope = -4/7